OpenStudy (anonymous):

Maria flips a dime four times. What is the probability that she will get at least two heads?

OpenStudy (anonymous):

2/8 or 1/4

OpenStudy (anonymous):

2/4 =1/2

OpenStudy (anonymous):

2/4=1/2

OpenStudy (anonymous):

I said 2/8 bc a coin has two sides qnd if you do it four times, that would be 8 and she gets two heads but im not sure since @lulu31302 and. @Jonathan101 both said 1/2

OpenStudy (anonymous):

@Heypips

OpenStudy (anonymous):

You flip 4 times but then you are calculating how many times it comes down heads. So flip and it lands on tales, flips lands on heads, flip lands on heads, flips lands on tails. Just did it, it is 2/4

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

but what would be the answer choice? @Jonathan101 @Heypips

OpenStudy (anonymous):

I'm confused now

OpenStudy (anonymous):

You could use the internet... Sorry @Jonathan101 confused me

OpenStudy (anonymous):

same

OpenStudy (anonymous):

11/16

OpenStudy (anonymous):

@n33dhelp1 sorry

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

did the math it is 11/16

OpenStudy (anonymous):

"There are 2^4 = 16 possible results given by 4 coin flips There are (4 C 2) = 6 configurations with exactly 2 heads, (4 C 3) = 4 with exactly 3 heads and (4 C 4) = 1 with exactly 4 heads. So probability (if coin is unbiased) is (6 + 4 + 1) / 16 = 11/16 (and probability that exactly 2 are heads is 3/8, not 1/2 as some people answer) " https://answers.yahoo.com/question/index?qid=20100207134141AAvPpXQ

OpenStudy (anonymous):

its ok @heypips